# Graph Coloring for Multi-Task Learning

### v2

@@ Abstract
When different objectives conflict with each other in multi-task learning, gradients begin to interfere and slow convergence, thereby potentially reducing the final model’s performance. To address this, we introduce SON-GOKU, a scheduler that computes gradient interference, constructs an interference graph, and then applies greedy graph-coloring to partition tasks into groups that align well with each other. At each training step, only one group (color class) of tasks are activated, and the grouping partition is constantly recomputed as task relationships evolve throughout training. By ensuring that each mini-batch contains only tasks that pull the model in the same direction, our method improves the effectiveness of any underlying multi-task learning optimizer without additional tuning. Since tasks within these groups will update in compatible directions, multi-task learning will improve model performance rather than impede it. Empirical results on six different datasets show that this interference-aware graph-coloring approach consistently outperforms baselines and state-of-the-art multi-task optimizers. We provide extensive theory showing why grouping and sequential updates improve multi-task learning, with guarantees on descent, convergence, and the ability to accurately identify what tasks conflict or align.
@@ 1 Introduction
Multi-task learning (MTL) trains a single model to solve several tasks simultaneously, sharing knowledge across them to learn more effectively (Caruana, 1997). This allows models to generalize better and converge faster. However, a key issue known as negative transfer arises when tasks don’t align very well with each other (Sener & Koltun, 2018; Shi et al., 2023). When two tasks push the shared network in different directions their gradients clash, slowing or even reversing learning. Prior work addresses this issue primarily via (1) gradient manipulation, which reshapes task gradients to reduce conflicts, and (2) loss reweighting, which rescales task objectives to balance their influence. While effective in some specific settings, these strategies typically treat conflict locally at the level of shared-parameter updates and often overlook the evolving global structure of interactions among tasks throughout training.
Some recent works focus on partitioning tasks into subsets (groups) and updating those groups separately. These approaches have been found to improve accuracy and training stability by forming groups with high measured affinity and then updating one group at a time (Fifty et al., 2021; Jeong & Yoon, 2025). Grouping can outperform gradient manipulation and loss reweighting when tasks form clusters with aligned gradients, because each update then reduces direct clashes in the shared layers, lowers gradient variance within the step, and lets compatible tasks reinforce one another while conflicting tasks wait for their turn.
However, grouping methods often face a few key limitations: (1) many rely on dense pairwise affinities that grow noisy and costly as the number of tasks rises (Fifty et al., 2021; Standley et al., 2020; Sherif et al., 2024), and (2) others predetermine or rarely update groups, so they drift as task relations change (Wang et al., 2024; Ruder, 2017), and (3) several use local heuristics that fail to enforce global compatibility or to specify how groups should rotate over time (Zhang & Yang, 2018; Malhotra et al., 2022).
@@ 2 Related Work
Many MTL methods (especially earlier ones) adjust task influence by learning or adapting loss weights. Examples include uncertainty-based scaling (Kendall et al., 2018), rate-based schemes such as DWA (Liu et al., 2019), and fast bilevel formulations like FAMO (Liu et al., 2023). FAMO in particular is notable for its $\mathcal{O}(1)$ per-step time complexity. These approaches keep all tasks active each step while modulating relative magnitudes. A completely different approach, which emerged in 2018 with MGDA (Sener & Koltun, 2018), focuses on updating shared-parameter update directions to mitigate interference. Methods like PCGrad (Yu et al., 2020), CAGrad (Liu et al., 2021), and MGDA (Sener & Koltun, 2018) modify the geometry of the shared update to reduce cross-task conflicts while still updating all tasks each step. A smaller body of work forms subsets of tasks to update together, using offline affinity estimation or training-dynamics signals (Fifty et al., 2021; Standley et al., 2020; Wang et al., 2024; Sherif et al., 2024). Most recently, Selective Task Group Updates proposes online grouping with sequential updates, reporting that update order can influence task-specific learning (Jeong & Yoon, 2025). SON-GOKU differs in mechanism from existing approaches (Section 4). It complements loss reweighting and gradient surgery, and we provide explicit guarantees on descent, convergence, and graph/partition recovery. An expanded discussion of more related work is provided in Appendix M.
@@ 3 Problem Setup
We formalize multi-task learning (MTL) (Caruana, 1997) as optimizing a shared network while activating only a subset of tasks at each step. Each task contributes a loss whose gradients may align or conflict. We quantify conflict using (the negative of) cosine similarity, embed tasks in a conflict graph, and later use that graph to derive a schedule. This section fixes notation and states the optimization goal that the proposed approach addresses.
@@ 4.1 Estimating Gradient Interference
which stabilizes cosine estimates while requiring only two buffers per task (current and previous). Whenever we refresh the schedule (every $R$ steps) we form the pairwise interference matrix
@@
Computing all K(K1)/2K(K-1)/2 cosines is $O(K^{2}d)$ with $d$ representing the shared‑parameter dimension. We also write $h_{k}^{(t)}=\nabla_{\phi_{k}}\mathcal{L}_{k}\!\bigl(\theta_{t},\phi_{k,t};\mathcal{B}_{k}^{(t)}\bigr)$ for the gradient with respect to the task-specific parameters $\phi_{k}$.
@@ 4.2 Conflict Graph Construction
To clarify, tasks are indexed by integers $1\ldots K$ in Equation 10. Edges connect tasks whose averaged gradients have cosine similarity less than $\tau$. Intuitively, larger $\tau$ yields a sparser conflict graph, typically fewer colors (larger per-step groups), and more frequent updates per task. Smaller $\tau$ results in a denser graph, more colors (smaller per-step groups), and less frequent updates per task.
@@ 4.3 Partitioning via Greedy Graph Coloring
We apply the Welsh-Powell largest-first greedy heuristic (Welsh & Powell, 1967) to color $G^{(r)}_{\tau}$ and obtain color classes $C^{(r)}_{1},\dots,C^{(r)}_{m_{r}}$. Classical graph-theory results (West, 2000; Diestel, 2017) guarantee the heuristic uses no more than $\Delta+1$ colors, where $\Delta$ is the maximum vertex degree. In practice $\Delta$ is small because many task pairs do not interfere, yielding concise schedules.
@@ 4.4.2 Warm-up and Annealing
We start with $\tau=1$ (no edges, full simultaneous training) for the first $T_{\mathrm{warm}}$ steps, then logarithmically anneal $\tau$ to a target value $\tau^{*}$. This mitigates noisy gradient signals early in training.
4.5 Time Complexity
The proposed scheduler has a time complexity of $\Theta(K^{2}d)$ per refresh. However, unlike many MTL approaches, our scheduler concentrates its extra work in occasional refreshes. This time complexity therefore becomes $\Theta\!\big(K^{2}d/R\big)$ amortized per training step where $R$ s the refresh period (the number of training steps between conflict-graph rebuilds). It adds non-trivial overhead which grows quadratically with $K$ (number of tasks) but shrinks as $R$ grows. We provide a full analysis of the time complexity in Appendix D and discuss approaches to reducing time complexity under certain conditions in Appendix D.5.
@@ 5 Theoretical Analysis
We discuss some of the main guarantees behind SON-GOKU. For a very brief overview: (1) Updating groups of tasks whose gradients are mostly low-conflict (no internal edges) reduces the objective on average and still achieves the usual $1/\sqrt{T}$ convergence rate; (2) Over a refresh window, scheduling several group updates can beat one mixed update that uses all tasks at once; and (3) With a small number of recent gradient measurements per task (via EMA) and a margin separating conflicts, the estimated conflict graph matches the ideal one, giving a short schedule where every task is updated at least once every $\Delta+1$ steps ($\Delta$ is the maximum number of conflicts for any task). We provide expanded assumptions, definitions, proofs, reasoning, analysis, etc. in Appendix 5.4I.
@@ 5.2 Nonconvex Convergence at the Standard Rate up to a Small Factor
When $\tau=0$, the constant matches the classical bound (Bottou et al., 2018; Ghadimi & Lan, 2013); as $\tau\to 1$, it at most doubles, matching the intuition that conflict can cancel up to half of the progress. This demonstrates that scheduling does not degrade asymptotic progress. SON-GOKU preserves the $1/\sqrt{T}$ decay of the gradient norm while controlling the constant through the compatibility threshold $\tau$. In other words, we keep the standard rate of SGD and trade a small constant for reduced interference.
@@ 5.3 When Scheduled Groups Outperform a Single Mixed Update
We compare two ways to use the same gradients gathered at a refresh: a scheduled sequence of per-group steps (i.e., the scheduler used in SON-GOKU) versus a single aggregated step. Using a telescoping $L$-smooth bound and evaluating both trajectories at a common linearization—i.e., expanding $F$ at the refresh start $\theta_{t_{r}}$ and applying the same first-order model with the same step size—the scheduled bound is never worse and is strictly better when cross-group interaction terms are sufficiently negative (so mixed updates would cancel progress).
@@ 5.4 Exact Recovery of the Population Conflict Graph and Task Partition
We show that, after observing gradients for only a modest number of steps, the scheduler can exactly reconstruct the true conflict relations among tasks by averaging recent gradients (EMA), computing pairwise cosines, thresholding at $\tau$, and coloring the resulting graph. Under a separation margin $\gamma$ around the threshold (tasks are meaningfully different), bounded noise, and bounded drift within each refresh window, the conflict graph estimated from finite data agrees, with high probability, with the ideal population conflict graph $G^{\star}\tau$ (defined from the pairwise cosines of the true mean gradients $\{\mu_{i}\}_{i=1}^{K}$ at the start of the refresh window). Equivalently, when the uniform cosine estimation error is below $\gamma$, we have $\widehat{G}\tau=G^{\star}\tau$ and the resulting grouping recovers the ground-truth task partition. This explains why the scheduler’s group structure is trustworthy and ties the required number of recent gradient measurements per task to interpretable quantities such as noise level, margin, and the number of tasks. For example, an effective sample size of $n_{\mathrm{eff}}\;\gtrsim\;\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\bigl(K/\delta\bigr)$ suffices in our analysis.
@@ 5.5 Scheduling Properties with Few Groups and Bounded Staleness
Welsh-Powell greedy coloring uses at most $\Delta+1$ colors on a graph whose maximum degree is $\Delta$ (Bonamy et al., 2018). Running the colors in a fixed cycle means each task is updated at least once every $m\leq\Delta+1$ steps. Equivalently, no task waits more than $\Delta$ steps between updates (bounded staleness).
This means that the schedule length is controlled by the worst conflict degree $\Delta$ rather than by the total number of tasks $K$. This results in two important benefits: (1) a minimum update-frequency guarantee, since every task receives an update at least once per cycle of length $\leq\Delta+1$; and (2) compatibility with standard bounded-delay conditions used in analyses of asynchronous SGD (e.g., Niu et al. 2011; Lian et al. 2015), with delay parameter at most $\Delta$. When $\Delta\ll K$, we achieve both low interference (few conflicts per step) and low staleness (short update gaps).
@@ 6.1 Datasets
We evaluate across six benchmarks spanning vision, multimodal, and time-series. For each dataset we specify a small set of primary tasks and add positive and negative auxiliaries to stress interference. Architectures are standard backbones (e.g., ResNet-18 for image tasks, CNN/BiLSTM for time-series) with task-specific heads. Full dataset and task definitions, auxiliary construction, and architecture details (including preprocessing and head designs) appear in Appendix J and Table 3.
@@ 6.3 Scheduler Extension Models
6.4 Ablation Study
6.4.1 Static One-Shot Coloring
We run the greedy graph coloring once at the start of training, freeze the resulting task groups, and never recompute the conflict graph. All other hyperparameters ($\tau$, history length $H$, and update interval $R$) match the full scheduler. As training progresses we expect the fixed coloring to grow stale, mixing tasks whose interference relationships have changed. This ablation isolates the benefit of dynamic recoloring, showing how much performance depends on adapting the schedule to evolving gradient conflicts.
@@ 7 Results and Discussion
Table 1: Performance of Evaluated Approaches Across Datasets
@@
CIFAR-10 AV‑MNIST MM‑IMDb Acc. (%) $\uparrow$ MAE $\downarrow$ Acc. (%) $\uparrow$ MAE $\downarrow$ Angle Error $\downarrow$ Seg. MIOU $\uparrow$ Depth RMSE $\downarrow$
Uniform 55 63 56 45 0.57 52 0.54 21.6 0.059 0.73
GradNorm 61 65 58 47 0.57 53 0.52 21.4 0.054 0.65
MGDA 59 62 56 44 0.57 53 0.53 21.8 0.63 0.75
PCGrad 61 65 58 50 0.55 58 0.48 20.9 0.07 0.69
CAGrad 59 62 57 46 0.58 53 0.52 21.9 0.65 0.73
AdaTask 63 67 59 47 0.59 55 0.52 20.3 0.69 0.65
FAMO 64 70 61 52 0.53 60 0.49 19.9 0.074 0.63
FairGrad 62 66 59 52 0.54 60 0.47 20.7 0.072 0.67
Nash-MTL 63 66 60 52 0.54 60 0.47 20.6 0.073 0.67
Static One-Shot 61 66 58 48 0.56 54 0.51 20.5 0.071 0.65
Single-Step 40 59 20 42 0.60 47 0.55 26.4 0.042 0.81
SON-GOKU ++ GradNorm 62 69 59 51 0.53 59 0.49 19.6 0.073 0.64
SON-GOKU ++ AdaTask 67 71 63 52 0.53 59 0.48 20.1 0.68 0.67
SON-GOKU ++ PCGrad 65 70 60 54 0.52 62 0.45 19.7 0.076 0.62
SON-GOKU 65 69 61 51 0.53 58 0.50 19.8 0.073 0.59
Results for all models across every experiment are depicted in Table 1. Across ten metrics on six datasets, our conflict-aware schedulers consistently match or exceed all baseline methods.

### v4

@@ Abstract
When different objectives conflict with each other in multi-task learning, gradients begin to interfere and slow convergence, thereby potentially reducing the final model’s performance. To address this, we introduce SON-GOKU, a scheduler that computes gradient interference, constructs an interference graph, and then applies greedy graph-coloring to partition tasks into groups that align well with each other. At each training step, only one group (color class) of tasks are activated, and the grouping partition is constantly recomputed as task relationships evolve throughout training. By ensuring that each mini-batch contains only tasks that pull the model in the same direction, our method improves the effectiveness of any underlying multi-task learning optimizer without additional tuning. Since tasks within these groups will update in compatible directions, multi-task learning will improve model performance rather than impede it. Empirical results on six different datasets show that this interference-aware graph-coloring approach consistently outperforms baselines and state-of-the-art multi-task optimizers. We provide extensive theory showing why grouping and sequential updates improve multi-task learning, with guarantees on descent, convergence, and the ability to accurately identify what tasks conflict or align.Code implementation is available at the following URL, note that it is temporarily anonymized to comply with double-blind peer review: https://anonymous.4open.science/r/SON-GOKU-ICLR-95AB
@@ 1 Introduction
Multi-task learning (MTL) trains a single model to solve several tasks simultaneously, sharing knowledge across them to learn more effectively (Caruana1997). This allows models to generalize better and converge faster. However, a key issue known as negative transfer arises when tasks don’t align very well with each other (Sener2018; Shi2023Recon). When two tasks push the shared network in different directions their gradients clash, slowing or even reversing learning. Prior work addresses this issue primarily via (1) gradient manipulation, which reshapes task gradients to reduce conflicts, and (2) loss reweighting, which rescales task objectives to balance their influence. While effective in some specific settings, these strategies typically treat conflict locally at the level of shared-parameter updates and often overlook the evolving global structure of interactions among tasks throughout training.
Some recent works focus on partitioning tasks into subsets (groups) and updating those groups separately. These approaches have been found to improve accuracy and training stability by forming groups with high measured affinity and then updating one group at a time (Fifty2021; jeong2025selective). Grouping can outperform gradient manipulation and loss reweighting when tasks form clusters with aligned gradients, because each update then reduces direct clashes in the shared layers, lowers gradient variance within the step, and lets compatible tasks reinforce one another while conflicting tasks wait for their turn.
However, grouping methods often face a few key limitations: (1) many rely on dense pairwise affinities that grow noisy and costly as the number of tasks rises (Fifty2021; Standley2020TaskGrouping; Sherif_2024); (2) others predetermine or rarely update groups, so they drift as task relations change (wang2024towards; Ruder2017Overview); and (3) several use local heuristics that fail to enforce global compatibility or to specify how groups should rotate over time (zhang2018overview; malhotra2022dropped).
@@ 2 Related Work
Prior work has identified the phenomenon of gradient interference in multi-task learning and explored several strategies to mitigate it. We group these such strategies into four families: (1) Tuned Loss Weighting, (2) Adaptive Loss Weighting, (3) Gradient-Level Conflict Mitigation, and (4) Empirical Task Grouping. SON-GOKU falls into family (4).
Many MTL methods (especially earlier ones) adjust task influence by learning or adapting loss weights. Examples include uncertainty-based scaling (kendall2018multi), rate-based schemes such as DWA (Liu2019EndToEnd), and fast bilevel formulations like FAMO (Liu2023FAMO). FAMO in particular is notable for its $\mathcal{O}(1)$ per-step time complexity. These approaches keep all tasks active each step while modulating relative magnitudes. A completely different approach, which emerged in 2018 with MGDA (Sener2018), focuses on updating shared-parameter update directions to mitigate interference. Methods like PCGrad (Yu2020PCGrad), CAGrad (Liu2021CAGrad), and MGDA (Sener2018) modify the geometry of the shared update to reduce cross-task conflicts while still updating all tasks each step. A smaller body of work forms subsets of tasks to update together, using offline affinity estimation or training-dynamics signals (Fifty2021; Standley2020TaskGrouping; wang2024towards; Sherif_2024). See Appendix Q for additional analysis of non-conflict task grouping. Most recently, Selective Task Group Updates proposes online grouping with sequential updates, reporting that update order can influence task-specific learning (jeong2025selective). SON-GOKU differs in mechanism from existing approaches (Section 4). It complements loss reweighting and gradient surgery, and we provide explicit guarantees on descent, convergence, and graph partition recovery. An expanded discussion and commentary of related work is provided in Appendix M.
@@ 3 Problem Setup
We formalize multi-task learning (MTL) (Caruana1997) as optimizing a shared network while activating only a subset of tasks at each step. Each task contributes a loss whose gradients may align or conflict. We quantify conflict using (the negative of) cosine similarity, embed tasks in a conflict graph, and later use that graph to derive a schedule (see Appendix P for a unique, modular, formulation and results with alternative measures of affinity). This section fixes notation and states the optimization goal that the proposed approach addresses.
@@ 4.1 Estimating Gradient Interference
which stabilizes cosine estimates while requiring only two buffers per task (current and previous). To estimate such cosines in practice, we employ low dimensional sketches of the EMA for each task, so the additional memory usage scales well (Ghashami2016). Whenever we refresh the schedule (every $R$ steps) we form the pairwise interference matrix.
@@
Computing all K(K1)/2K(K-1)/2 cosines via the Gram matrix in the $r$-dimensional sketch space costs $O(Kdr+K^{2}r),$ namely $O(Kdr)$ to form the sketch $M_{f}$ and $O(K^{2}r)$ for the Gram product, where $r\ll d$ is the sketch width (see Appendix D.5.1). We also write $h_{k}^{(t)}=\nabla_{\phi_{k}}\mathcal{L}_{k}\!\bigl(\theta_{t},\phi_{k,t};\mathcal{B}_{k}^{(t)}\bigr)$ for the gradient with respect to the task-specific parameters $\phi_{k}$.
@@ 4.2 Conflict Graph Construction
To clarify, tasks are indexed by integers $1\ldots K$ in Equation 10. Edges connect tasks whose averaged gradients have cosine similarity less than $\tau$. Intuitively, larger $\tau$ yields a sparser conflict graph, typically fewer colors (larger per-step groups), and more frequent updates per task. Smaller $\tau$ results in a denser graph, more colors (smaller per-step groups), and less frequent updates per task. This construction reflects optimization-time interference. $G^{(r)}_{\tau}$ is symmetric and undirected, derived from current gradient geometry to decide which tasks should not be updated together.
@@ 4.3 Partitioning via Greedy Graph Coloring
We apply the Welsh-Powell largest-first greedy heuristic (Welsh1967) to color $G^{(r)}_{\tau}$ and obtain color classes $C^{(r)}_{1},\dots,C^{(r)}_{m_{r}}$. Classical graph-theory results (West2000; Diestel2017) guarantee the heuristic uses no more than $\Delta+1$ colors, where $\Delta$ is the maximum vertex degree. In practice $\Delta$ is small because many task pairs do not interfere, yielding concise schedules.
@@ 4.4.2 Warm-up and Annealing
We start with $\tau=1$ (no edges, full simultaneous training) for the first $T_{\mathrm{warm}}$ steps, then logarithmically anneal $\tau$ to a target value $\tau^{*}$. This mitigates noisy gradient signals early in training. Similarly, we can set the refresh period with a smaller $R$ to adapt to changing gradients and increase it as training stabilizes (Appendix O).
4.5 Time Complexity and Space Complexity
Using the sketched implementation described in Appendix D, a single refresh of the SON-GOKU scheduler has time complexity $O(Kdr+K^{2}r),$ where $r\ll d$ is the sketch width. However, unlike many MTL approaches, our scheduler concentrates its extra work in occasional refreshes. This time complexity therefore becomes $O\!\left(\frac{Kr(d+K)}{R}\right)$ amortized per training step where $R$ is the refresh period (the number of training steps between conflict-graph rebuilds).For the small, fixed $r$ used in our experiments, this overhead still grows roughly quadratically in $K$ but is independent of $d$ up to the $O(Kdr)$ sketching term and shrinks linearly with the refresh period $R$. Similarly, SON-GOKU’s persistent space complexity of $O(K^{2})$ scales with $K$ but not $d$, the number of model parameter dimensions, allowing it to maintain low memory usage even with large backbone models. We provide a full analysis of the time complexity in Appendix D and discuss approaches to reducing time complexity under certain conditions in Appendix D.5. See also Appendix R for scaling behavior with larger backbones.
@@ 5 Theoretical Analysis
We discuss some of the main guarantees behind SON-GOKU. For a very brief overview: (1) Updating groups of tasks whose gradients are mostly low-conflict (no internal edges) reduces the objective on average and still achieves the usual $1/\sqrt{T}$ convergence rate; (2) Over a refresh window, scheduling several group updates can beat one mixed update that uses all tasks at once; and (3) With a small number of recent gradient measurements per task (via EMA) and a margin separating conflicts, the estimated conflict graph matches the ideal one, giving a short schedule where every task is updated at least once every $\Delta+1$ steps ($\Delta$ is the maximum number of conflicts for any task). We provide expanded assumptions, definitions, proofs, reasoning, analysis, etc. in Appendix 5.4I (see also Appendix N, PR).
@@ 5.2 Nonconvex Convergence at the Standard Rate up to a Small Factor
When $\tau=0$, the constant matches the classical bound (bottou2018optimization; GhadimiLan2013); as $\tau\to 1$, it at most doubles, matching the intuition that conflict can cancel up to half of the progress. This demonstrates that scheduling does not degrade asymptotic progress. SON-GOKU preserves the $1/\sqrt{T}$ decay of the gradient norm while controlling the constant through the compatibility threshold $\tau$. In other words, we keep the standard rate of SGD and trade a small constant for reduced interference.
@@ 5.3 When Scheduled Groups Outperform a Single Mixed Update
We compare two ways to use the same gradients gathered at a refresh: a scheduled sequence of per-group steps (i.e., the scheduler used in SON-GOKU) versus a single aggregated step. Using a telescoping $L$-smooth bound and evaluating both trajectories at a common linearization (i.e., expanding $F$ at the refresh start $\theta_{t_{r}}$ and applying the same first-order model with the same step size) the scheduled bound is never worse and is strictly better when cross-group interaction terms are sufficiently negative (so mixed updates would cancel progress).
@@ 5.4 Exact Recovery of the Population Conflict Graph and Task Partition
We show that, after observing gradients for only a modest number of steps, the scheduler can exactly reconstruct the true conflict relations among tasks by averaging recent gradients (EMA), computing pairwise cosines, thresholding at $\tau$, and coloring the resulting graph. Under a separation margin $\gamma$ around the threshold (tasks are meaningfully different), bounded noise, and bounded drift within each refresh window, the conflict graph estimated from finite data agrees, with high probability, with the ideal population conflict graph $G^{\star}\tau$ (defined from the pairwise cosines of the true mean gradients $\{\mu_{i}\}_{i=1}^{K}$ at the start of the refresh window). Equivalently, when the uniform cosine estimation error is below $\gamma$, we have $\widehat{G}_{\tau}=G^{\star}_{\tau}$ and the resulting grouping recovers the ground-truth task partition. This explains why the scheduler’s group structure is trustworthy and ties the required number of recent gradient measurements per task to interpretable quantities such as noise level, margin, and the number of tasks. For example, an effective sample size of $n_{\mathrm{eff}}\;\gtrsim\;\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\bigl(K/\delta\bigr)$ suffices in our analysis.
@@ 5.5 Scheduling Properties with Few Groups and Bounded Staleness
Welsh-Powell greedy coloring uses at most $\Delta+1$ colors on a graph whose maximum degree is $\Delta$ (bonamy2018boundingchifractiondelta). Running the colors in a fixed cycle means each task is updated at least once every $m\leq\Delta+1$ steps. Equivalently, no task waits more than $\Delta$ steps between updates (bounded staleness).
This means that the schedule length is controlled by the worst conflict degree $\Delta$ rather than by the total number of tasks $K$. This results in two important benefits: (1) a minimum update-frequency guarantee, since every task receives an update at least once per cycle of length $\leq\Delta+1$; and (2) compatibility with standard bounded-delay conditions used in analyses of asynchronous SGD (e.g., Recht2011Hogwild; lian2015asynchronous), with delay parameter at most $\Delta$. When $\Delta\ll K$, we achieve both low interference (few conflicts per step) and low staleness (short update gaps).
@@ 6.1 Datasets
We evaluate across six benchmarks spanning vision, multimodal, and time-series. For each dataset we specify a small set of primary tasks and add positive and negative auxiliaries to stress interference. Architectures are standard backbones (e.g., ResNet-18 for image tasks, CNN/BiLSTM for time-series) with task-specific heads. Full dataset and task definitions, auxiliary construction, and architecture details (including preprocessing and head designs) appear in Appendix J and Table 4. We provide additional experiments with varying backbones in Appendix R.
@@ 7 Results and Discussion
Table 1: Performance of Evaluated Approaches Across Datasets. DM represents Density-Matched ablation variants
@@
CIFAR-10 AV-MNIST $M$-IMDb Acc. (%) $\uparrow$ MAE $\downarrow$ Acc. (%) $\uparrow$ MAE $\downarrow$ Angle Error $\downarrow$ Seg. MIOU $\uparrow$ Depth RMSE $\downarrow$
Uniform $55$ $\pm 2.2$ $63$ $\pm 1.5$ $56$ $\pm 2.8$ $45$ $\pm 2.4$ $0.57$ $\pm 0.030$ $52$ $\pm 2.0$ $0.54$ $\pm 0.024$ $21.6$ $\pm 0.27$ $0.059$ $\pm 0.003$ $0.73$ $\pm 0.018$
GradNorm $61$ $\pm 1.6$ $65$ $\pm 1.1$ $58$ $\pm 2.0$ $47$ $\pm 2.3$ $0.57$ $\pm 0.020$ $53$ $\pm 2.1$ $0.52$ $\pm 0.019$ $21.4$ $\pm 0.23$ $0.054$ $\pm 0.004$ $0.65$ $\pm 0.016$
MGDA $59$ $\pm 2.9$ $62$ $\pm 1.7$ $56$ $\pm 3.3$ $44$ $\pm 3.0$ $0.57$ $\pm 0.036$ $53$ $\pm 2.5$ $0.53$ $\pm 0.030$ $21.8$ $\pm 0.33$ $0.063$ $\pm 0.005$ $0.75$ $\pm 0.024$
PCGrad $61$ $\pm 1.9$ $65$ $\pm 1.3$ $58$ $\pm 2.3$ $50$ $\pm 2.1$ $0.55$ $\pm 0.024$ $58$ $\pm 2.0$ $0.48$ $\pm 0.021$ $20.9$ $\pm 0.24$ $0.070$ $\pm 0.004$ $0.69$ $\pm 0.013$
CAGrad $59$ $\pm 2.0$ $62$ $\pm 1.1$ $57$ $\pm 2.5$ $46$ $\pm 2.5$ $0.58$ $\pm 0.031$ $53$ $\pm 1.9$ $0.52$ $\pm 0.024$ $21.9$ $\pm 0.29$ $0.065$ $\pm 0.004$ $0.73$ $\pm 0.018$
AdaTask $63$ $\pm 1.5$ $67$ $\pm 0.9$ $59$ $\pm 1.9$ $47$ $\pm 1.9$ $0.59$ $\pm 0.026$ $55$ $\pm 2.2$ $0.52$ $\pm 0.024$ $20.3$ $\pm 0.23$ $0.069$ $\pm 0.004$ $0.65$ $\pm 0.015$
FAMO $64$ $\pm 1.2$ $70$ $\pm 1.0$ $61$ $\pm 1.6$ $52$ $\pm 2.0$ $0.53$ $\pm 0.021$ $60$ $\pm 1.8$ $0.49$ $\pm 0.018$ $19.9$ $\pm 0.19$ $0.074$ $\pm 0.003$ $0.63$ $\pm 0.012$
FairGrad $62$ $\pm 1.8$ $66$ $\pm 1.3$ $59$ $\pm 2.5$ $52$ $\pm 2.5$ $0.54$ $\pm 0.026$ $60$ $\pm 2.0$ $0.47$ $\pm 0.022$ $20.7$ $\pm 0.27$ $0.072$ $\pm 0.004$ $0.67$ $\pm 0.015$
Nash-MTL $63$ $\pm 1.9$ $66$ $\pm 1.2$ $60$ $\pm 2.1$ $52$ $\pm 2.3$ $0.54$ $\pm 0.024$ $60$ $\pm 2.3$ $0.47$ $\pm 0.023$ $20.6$ $\pm 0.24$ $0.073$ $\pm 0.004$ $0.67$ $\pm 0.013$
Static One-Shot $61$ $\pm 2.0$ $66$ $\pm 1.1$ $58$ $\pm 2.6$ $48$ $\pm 2.3$ $0.56$ $\pm 0.027$ $54$ $\pm 2.1$ $0.51$ $\pm 0.025$ $20.5$ $\pm 0.25$ $0.071$ $\pm 0.004$ $0.65$ $\pm 0.016$
Single-Step $40$ $\pm 4.2$ $59$ $\pm 2.4$ 2020 $\pm 5.4$ $42$ $\pm 3.9$ $0.60$ $\pm 0.041$ $47$ $\pm 3.5$ $0.55$ $\pm 0.034$ $26.4$ $\pm 0.55$ $0.042$ $\pm 0.006$ $0.81$ $\pm 0.029$
SON-GOKU (Threshold, DM) 63 68 59 49 0.55 56 0.51 20.6 0.071 0.61
SON-GOKU (kNN-Symm.) 60 65 55 46 0.57 52 0.53 22.1 0.066 0.70
SON-GOKU (kNN-Symm., DM) 61 66 57 47 0.56 54 0.52 21.4 0.068 0.66
SON-GOKU (Signed-only) 56 63 52 43 0.60 50 0.56 24.0 0.053 0.76
SON-GOKU (Signed-only, DM) 58 64 54 45 0.59 52 0.54 23.0 0.056 0.73
SON-GOKU (Quantile) 64 68 60 50 0.54 57 0.50 20.3 0.072 0.60
SON-GOKU (Quantile, DM) 65 69 61 51 0.53 58 0.50 20.0 0.072 0.59
SON-GOKU ++ GradNorm $62$ $\pm 1.4$ $69$ $\pm 1.0$ $59$ $\pm 1.7$ $51$ $\pm 1.8$ $0.53$ $\pm 0.022$ $59$ $\pm 1.7$ $0.49$ $\pm 0.018$ $\mathbf{19.6}$ $\pm 0.19$ $0.073$ $\pm 0.003$ $0.64$ $\pm 0.011$
SON-GOKU ++ AdaTask $\mathbf{67}$ $\pm 1.2$ $\mathbf{71}$ $\pm 0.9$ $\mathbf{63}$ $\pm 1.6$ $52$ $\pm 1.7$ $0.53$ $\pm 0.021$ $59$ $\pm 1.8$ $0.48$ $\pm 0.017$ $20.1$ $\pm 0.20$ $0.068$ $\pm 0.004$ $0.67$ $\pm 0.013$
SON-GOKU ++ PCGrad $65$ $\pm 1.3$ $70$ $\pm 0.9$ $60$ $\pm 1.8$ $\mathbf{54}$ $\pm 2.0$ $\mathbf{0.52}$ $\pm 0.024$ $\mathbf{62}$ $\pm 1.6$ $\mathbf{0.45}$ $\pm 0.020$ $19.7$ $\pm 0.18$ $\mathbf{0.076}$ $\pm 0.003$ $0.62$ $\pm 0.010$
SON-GOKU $65$ $\pm 1.5$ $69$ $\pm 1.0$ $61$ $\pm 1.8$ $51$ $\pm 1.9$ $0.53$ $\pm 0.023$ $58$ $\pm 1.7$ $0.50$ $\pm 0.018$ $19.8$ $\pm 0.20$ $0.073$ $\pm 0.004$ $\mathbf{0.59}$ $\pm 0.012$
Results for all models across every experiment are depicted in Table 1. All metrics are held-out test results under identical training setups and architectures. Across ten metrics on six datasets, our conflict-aware schedulers consistently match or exceed all baseline methods.